OpenStudy (anonymous):

HELP HELP HELP SOS SOS SOS !!! HELP WANTED !!! REWARD:MEDAL AND NEW FAN! PLZ PLZ PLZ PLZ PLZ PLZ PLZ PLZ HELP!!!

OpenStudy (anonymous):

what

OpenStudy (anonymous):

whts ur question

OpenStudy (anonymous):

6 * 94 = 6 * 9 + 6 * 4 = 78

OpenStudy (anonymous):

THANKS AND THE 2ND ?

OpenStudy (anonymous):

P = 2L + 2W = 2 * 15.3 + 2 * 2.1 = 2 * (15.3 + 2.1) = 34.8

OpenStudy (anonymous):

seriously thats easy

OpenStudy (anonymous):

THX :)